Suppose a consumer’s utility function is given a U = 100X0.25Y0.75.The prices of the two commodities X and Y are Birr 2 and Birr 5 per unit respectively. If the consumer’s income is Birr 280, how many units of each commodity should the consumer buy to maximize his/her utility? 


1
Expert's answer
2021-08-05T18:02:32-0400

U = 100X0.25Y0.75

Budget constraint is

"2x + 5y = 280"

Maximize utility function

subject to the budget constraint

Max U = 100X0.25Y0.75

"2x + 5y \u2264 280"

"x\u22650 ,\ny\u22650"


We introduce a lagrangian function to our function

"L =" 100X0.25Y0.75 + ƛ ("280- 2x - 5y" )

FIRST ORDER CONDITIONS

ɚL/ɚX = 25X-0.75Y0.75 - ƛ2 = 0 .......EQUATION 1

ɚL/ɚY=75X0.25Y-0.25- ƛ5 = 0 ......... EQUATION 2

ɚL/ɚ"\u019b =" ƛ = 280- 2x - 5y = 0 ..............EQUATION 3

WE COMBINE EQUATION 1 AND 2

 ƛ=(25X-0.75Y0.75 ) ÷ 2 = (75X0.25Y-0.25 ) ÷ 5

SOLVING THIS WE GET;

5 ÷6 =X ÷ Y

Thus x= 0.833y

y= 1.2x

Using equation three we substitute one of the above rquations

280- 2x - 5y = 0

2x- 5y=280 but x= 0.833y

0.833y x 2 - 5y=280

Solving this we get

y= 168 units

From y= 1.2x we substitute y = 168 and solve to find x

168 = 1.2x

x = 140 units

Therefore the consumer maximizes utility at y= 168 units and x = 140 units